OpenStudy (anonymous):

My conjecture is n^3 - n is divisible by 6. I have done the basis step, and now have to do the inductive step, and i'm kinda stuck in that. Help??

hartnn (hartnn):

so now you want to prove that n(n+1)(n+2) is divisible by 6 , right ??

OpenStudy (aravindg):

well did u write equation in k ?

OpenStudy (anonymous):

yes, and yes :P

OpenStudy (shubhamsrg):

@hartnn n(n+1)(n-1) *

hartnn (hartnn):

take 2 cases, n odd n even

hartnn (hartnn):

whats that shubham ?

OpenStudy (anonymous):

this is where i'm at right now: \[k^3 -6m(k+1)(k+2)\]

hartnn (hartnn):

hmm...how ? :P

OpenStudy (shubhamsrg):

n^3 - n = n(n^2 -1) = n(n+1)(n-1) how did you write n(n+1)(n+2) ?

hartnn (hartnn):

i thought using induction, u assume the result for 'n' and prove it for 'n+1'

OpenStudy (aravindg):

@tanvidais13 can u show hw u gt to that?

OpenStudy (anonymous):

because k was outside, and so then I replaced it with k^3 + 6m

hartnn (hartnn):

the case when we have to show the result for 'n+1' (n+1)^3-(n+1)= n(n+1)(n+2)

OpenStudy (anonymous):

because k^3 - k = 6m, where m is any integer (showing that it is divisible by 6)

OpenStudy (anonymous):

ideally, i would have liked to get 6m(something), and therefore it is divisble by 6. Right??

OpenStudy (anonymous):

OH WAIT OMG I JUST GOT IT

hartnn (hartnn):

got it ? good :P

OpenStudy (aravindg):

yep infact use k^3 = 6m+k

OpenStudy (anonymous):

no.

OpenStudy (aravindg):

then?

OpenStudy (anonymous):

because i got the k(k+1)(k+2) thing, but how do I show that its divisble by 6??

hartnn (hartnn):

i didn't know my presence have that effect! ppl get answers on their own.....

OpenStudy (aravindg):

lols @hartnn

hartnn (hartnn):

ohh... take 2 cases k even k odd

OpenStudy (anonymous):

oh but lulz.

OpenStudy (anonymous):

but i can't do that. I have to 'induct' *nods head wisely*

hartnn (hartnn):

when k= odd, k+1 is even and either k or k+2 is divisible by 3.

hartnn (hartnn):

you already used induction and now you want to prove k(k+1)(k+2) is divisible by 6, anyhow.....

OpenStudy (anonymous):

^yup.

hartnn (hartnn):

when k= even, k+1 is odd and either k+1 or k+2 is divisible by 3. hence in either case....

OpenStudy (aravindg):

i just got something ,well we have \[k(k+1)(k-1)=6m\] we hav to prove \[k(k+1)(k+2) \text{ is divisible by 6 }\] then why dont u substitute \[k(k+1)\] from first equation?

OpenStudy (aravindg):

isnt that direct method ?

OpenStudy (anonymous):

wait i don't get it.

OpenStudy (aravindg):

@hartnn what do you say?

OpenStudy (anonymous):

but i've already substituted k^3 for 6m + k

OpenStudy (aravindg):

@tanvidais13 mine is a fresh attempt by me dont confuse with ur working

OpenStudy (anonymous):

oh ok.

OpenStudy (aravindg):

did u understand it now @tanvidais13 ?

OpenStudy (anonymous):

but that means dividing. and that's painful.

OpenStudy (blurbendy):

k(k+1)(k+2) = k(k^2 + 3k + 2) k^3 + 3k^2 + 2k 3(1/3k^3 + k^2 + 2/3k) Since this is in the form of 3(some stuff), and 6 is a multiple of 3, the k(k+1)(k+2) is divisible by 6. I don't know if this works, just thought I'd try.

OpenStudy (blurbendy):

then*

OpenStudy (aravindg):

@blurbendy by your metjod i can make any expression divisible by 6 it seems :P

OpenStudy (anonymous):

ok, i think i've got the answer. It's a mixture of mine and Arvind's method, but ok :P Anyhoo, thanks guyss :D

OpenStudy (aravindg):

:) seems ?

OpenStudy (anonymous):

and ok, stay on openstudy for a while, 'cuz there are going to be some more questions coming your way,

OpenStudy (aravindg):

yw ... :) post them separately ..we will try :)

OpenStudy (anonymous):

ok yay :D

OpenStudy (aravindg):

TIME for me and pikachu to carry on with our adventures lols